How many different committees of three students can be formed from a class of five students? Students can be on more than one committee.

Answer by stanbon(75887)

How many different committees of three students can be formed from a class of five students? Students can be on more than one committee.
-------------------------
Ans: 5C3 = 5!/[(5-3)!*3!] = (5*4)/(1*2) = 10
------------------------------------------------
Cheers,
Stan H.